Mussels are Canada’s top shellfish aquaculture product, produced in Newfoundland and Labrador, Nova Scotia, Prince Edward Island and Quebec. Mussels filter the sea water between their shells and accumulate viral, bacterial, and parasitic pathogens in their tissues. Among the mussels collected from a certain Canadian coastline 80% are infected with a parasite. A biologist examines 120 randomly selected mussels in this region. What is the probability that more than 86% of the mussels in this sample are infected with parasite?
Answer #1

Solution

\sigma\hat p =  \sqrt[p( 1 - p ) / n] = \sqrt [(0.80 * 0.20) / 120 ] = 0.0365

P( \hat p > 0.86) = 1 - P( \hat p < 0.86)

= 1 - P(( \hat p - \mu \hat p ) / \sigma \hat p < (0.86 - 0.80) / 0.0365)

= 1 - P(z < 1.64)

= 0.0505

Probability = 0.0505